If for a certain BCC crystal, the (110) plane has a separation of 1.181A and wavelength of 1.540A



Find the glance angle for the 1st line of the reflection.



B. Hence find the atomic radius

1
Expert's answer
2022-01-23T22:16:02-0500

Bragg's Law 

"n\lambda= 2dSin\theta"

glance angle for the 1st line of the reflection.


"\theta = Sin^{-1}(\frac{1\times 1.540}{2\times 1.181}) =40.69^0"


atomic radius

"R = \frac{\sqrt3}{4}a= \frac{\sqrt3}{4}1.181 = 0.511"